### v4

@@ 1 Introduction
Edge detection is a fundamental problem in computer vision and image processing, playing a crucial role in a wide range of applications, from autonomous driving to medical imaging. It serves as the cornerstone for many high-level computer vision tasks by identifying the boundaries and structures within images, such as salient object detection [1, 2, 3, 4, 5, 6], objection recognition [7, 8, 9, 10, 11], image segmentation [12, 13, 14, 15, 16, 17, 18, 19, 20], and object tracking [21, 22, 23]. With the rapid development of deep learning techniques, numerous researchers aim to employ Deep Convolutional Neural Networks (DCNNs) for edge detection. Recent studies have proposed several remarkable DCNN-based methods, such as HED [24], RCF [25], and CHRNet [26]. These DCNN-based approaches have achieved state-of-the-art performance on a number of benchmarks including BSDS500 [27] and NYUD-V2 [28], which demonstrate the strong power of DCNN.
@@
In this work, we attempt to tackle this central issue by addressing the two subsidiary problems. Our goal is to make the network generate one-pixel width edge lines without using the post-processing operation. On the one hand, most DCNN-based methods [24, 25, 29, 30, 31] leverage the powerful feature extraction capabilities of deep learning for edge detection. However, these methods neglect the utilization of prior knowledge regarding image edges, which are typically determined by abrupt changes in pixels. Image second-order derivative captures such changes and it precisely locates the positions of edge pixels by the zero-crossing points. Therefore, leveraging this characteristic can help DCNNs to locate the true edge pixels more accurately. In addition, the mere introduction of second-order derivative information can result in noisy edge maps, as illustrated in Fig. 1 (c). In response to this circumstance, we combine the second-order derivative information with the multi-scale contextual information, developing a Second-order Derivative-based Multi-scale Contextual Enhancement Module (SDMCM). The SDMCM is based on the traditional Laplacian operator [32] which can provide image second-order derivative cues, and dilated convolution with a large receptive field to capture the long-range multi-scale contextual information which can filter out the noise. The proposed SDMCM empowers the network to attain higher precision in true edge pixel localization, consequently generating crisp and clean edge maps.
@@ 2 Related work
Traditional methods: Early edge detection methods typically calculate image derivatives to produce edge maps. The Roberts operator [40] was a simple first-order derivative operator that detected edges by computing differences between diagonally adjacent pixels. The Sobel operator [41] was another first-order edge detector that used two $3\times 3$ kernels to calculate image gradients in horizontal and vertical directions. The Canny detector [42] was a robust algorithm that detected edges by reducing noise, calculating gradients, and applying non-maximum suppression and hysteresis thresholding. The Laplacian detector [32] located edge pixels by computing the second-order derivative of the image intensity, highlighting regions of rapid intensity change. Over time, some researchers improved edge detection by integrating texture, gradient, and other low-level features. Methods like Pb [43], gPb [27], and SE [44] used a classifier to generate object-level boundaries by utilizing these features. Despite their enhanced performance over derivative-based methods, they still relied on human-designed features and lacked semantic information, limiting further improvement.
@@
Deep learning-based methods: Deep learning-based techniques significantly advanced the area of edge detection and played a crucial role in most state-of-the-art (SOTA) edge detection methods. SOTA edge detectors in recent years have mainly adopted convolutional neural networks. These methods achieved significant performance with higher F-scores and some of them even surpassed humans on benchmarks such as BSDS500 dataset. Xie et al. [24] presented the first end-to-end edge detection architecture named Holistically-nested Edge Detection (HED), which was built on a fully convolutional VGG-16 network. They generated edge maps by fusing five side-output features with different scales and constructed a weighted cross-entropy loss function to address the problem of imbalanced pixel distribution. Based on HED, Liu et al. [25] further utilized multi-scale features by aggregating all features from different convolutional layers in each stage of VGG-16. This improved the network’s ability to capture contextual information, making it the first to outperform humans on BSDS500. He et al. [29] proposed an innovative bi-directional cascade network architecture, which leveraged information flow from shallow-to-deep and deep-to-shallow to capture multi-scale features within images comprehensively, significantly enhancing the performance. Soria et al. [45] proposed DexiNed, a network inspired by HED and Xception [46], which could produce detailed edge maps that were visually appealing without any pre-training or fine-tuning process. Su et al. [33] provided a lightweight yet effective solution for edge detection by integrating traditional edge detection operators into vanilla convolution in modern DCNN.
@@
Regarding precise edge detection, various methods have contributed significantly to the field. Wang et al. [47] focused on guiding the network to predict sharp edge maps by employing sub-pixel convolution to upsample features. Deng et al. [48] explained the reason for edge thickness and proposed a new loss function based on the Dice coefficient [49] which enables the network to generate crisp boundaries without requiring post-processing. Deng et al. [50] proposed DSCD to produce high-quality edge maps by introducing a novel loss function inspired by SSIM [51] and built a dense connection [52] hyper-module based on dilated convolution [53]. Cao et al. [30] presented an innovative network by stacking refinement modules and implementing an adaptive weighting strategy for the optimal combination between different loss functions, resulting in a competitive performance. Huan et al. [54] presented a context-aware tracing strategy using a novel tracing loss and a context-aware fusion block for crisp edge detection. Xuan et al. [55] proposed a novel network architecture designed to enhance the accuracy of edge detection through fine-scale corrective learning mechanisms. Liu et al. [56] proposed GCB-Net, a U-shape encoder-decoder architecture with feature enhancement and fusion modules, which could generate crisp and clean edge maps through multi-scale features and a novel mixed loss function based on Tversky index. Pu et al. [57] EDTER, the first Transformer-based [58] edge detector with a two-stage architecture. The model could extract well-defined boundaries of objects and meaningful edges by utilizing contextual information from the entire image, as well as detail local cues. Zhou et al. [59] designed an edge granularity network to estimate edge granularity from annotations and integrated it into both spatial and frequency domains of feature maps, enabling the generation of diverse edge predictions ranging from coarse object contours to fine textures. Cetinkaya et al. [60] constructed a ranking-based approach to address both class imbalance and label uncertainty in edge detection through ranking positive pixels over negatives and promoting high confidence edge pixels.
@@
Despite their remarkable performance, existing methods still fall short in adequately addressing the challenge of edge thickness. Our approach is motivated by these pioneering works [48, 34, 32, 56] to address the problem of edge thickness. Through the combination of multi-scale contextual and second-order derivative information, we enable the network to produce crisp and clean edge maps. Additionally, our method successfully addresses the issue of imbalanced pixel distribution by constructing a hybrid focal loss function, resulting in excellent performance.
@@ 4.1 Implementation details
Our network is built using the Pytorch deep learning framework [63]. During the training phase, the hyperparameters are as follows: the mini-batch size is 8, the initial learning rate is $1\times 10^{-4}$, the learning rate decay is 0.1, the weight decay is $5\times 10^{-4}$ and the number of training epochs is 40. We decay the learning rate every 5 epochs and adopt the Adam [64] method for optimization. All the experiments are performed using a single Tesla A40 GPU.
@@ 4.3 Ablation experiments on BSDS500 dataset
Loss function: Firstly, we perform ablation experiments on various loss functions to demonstrate the effectiveness of our HFL. To identify the optimal configuration of HFL, we conduct a series of comparison experiments with different parameters, and the comparison results are shown in rows 1-6 of Table 1. It can be observed that, as $\lambda$ decreases from 1 to 0.001, there is a notable increase in C-Eval metrics, while S-Eval metrics maintain relatively stable performance. Such observation can be explained by the complementary roles of the two loss components: focal Tversky loss serves to regulate the crispness of edge delineation, whereas focal loss ensures the accuracy of pixel-level classification in the predicted edge maps. Therefore, the optimal value of $\lambda$ is 0.001. Additionally, when fixed $\lambda=0.001$, the best results of both S-Eval and C-Eval are obtained at $\beta=0.7$, outperforming other values (0.6 and 0.8). In rows 11-14, it can be observed that the best performance is obtained at $\gamma=0.75$. Furthermore, we compare our HFL loss with other loss functions, including Weighted Cross-Entropy [24] (WCE), Hybrid Quality Focal Loss [65] (HQFL), and Hybrid Varifocal Loss [66] (HVFL). The comparison results shown in rows 7-10 demonstrate that the HFL maintains competitive performance in S-Eval (ODS=0.805, OIS=0.827, AP=0.805). More importantly, HFL demonstrates superior performance across all metrics in C-Eval (ODS=0.698, OIS=0.705, AP=0.754), notably outperforming WCE, HQFL, and HVFL. These results strongly indicate that our proposed HFL effectively enhances the model’s capability to produce crisp and clean edge maps while maintaining robust performance in pixel-level classification.
@@
Compression ratio: We also explore the effect of different channel compression ratios $r$ in SDMCM on the model performance and the experiment results are shown in rows 11-13 of Table 1 and Fig. 7. As the ratio increases, the number of model parameters also increases, resulting in a higher computational cost. However, the performance does not improve with an increasing compression ratio. The best performance is obtained at $r=\frac{1}{4}$. The reason for this phenomenon is that a higher compression ratio causes overfitting, resulting in a lower performance both in S-Eval and C-Eval. Therefore, we set the compression ratio to $\frac{1}{4}$, which achieves a balance between performance and parameters.
@@
Image derivative information: In addition, we investigate the effectiveness of different derivative operators by replacing the Laplacian operator in SDMCM with first-order derivative operators (Sobel and Scharr). As shown in rows 14-16 of Table 1, both alternative operators lead to consistent performance degradation across all metrics, substantiating our theoretical analysis that second-order derivative information provides more discriminative features for edge detection compared to first-order derivatives.
@@ 4.4 Comparison with SOTA methods
BSDS500: Firstly, we compare our method with top-performing algorithms on BSDS500. We select recent state-of-the-art edge detectors which can be divided into two categories: the first category is methods without deep learning, which includes Canny [42], gPb-UCM [27], SE [44], SCG [67], PMI [68], OEF [69], MES [70], and Liu et al. [71]; the second category is approaches using deep learning technique, which including DeepContour [72], DeepEdge [73], HED [24], RCF [25], BDCN [29], CED [47], LPCB [48], DRC [30], PiDiNet [33], CATS [54], FCL-Net [55], EDTER [57], Boosting-ED [74], CHRNet [26], DexiNed [45], BetaNet [75], GCB-Net [56], EdgeSAM [76], MuGE [59], and RankED [60]. We additionally reference the studies of prior researchers who utilize extra training data sourced from the PASCAL VOC Context dataset [77] and adopt multi-scale testing to further improve the performance. Table 3 presents all metrics including S-Eval, C-Eval, Params, and FPS. Fig. 8 shows some qualitative comparison results, and Fig. 9 draws the Precision-Recall curves in S-Eval.
@@
NYUD-V2: Secondly, we select the NYUD-V2 dataset to conduct another set of comparison experiments. We adopt some methods as before, which consists of algorithms without using deep learning such as gPb-UCM [27], OEF [69], gPb+NG [78], SE [44] and SE+NG+ [79], and recent top edge detectors based on deep learning such as HED [24], RCF [25], BDCN [29], CATS [54], DRC [30], PiDiNet [33], DexiNed [45], LPCB [48], CHRNet [26], BetaNet [75], GCB-Net [56], EDTER [57], Boosting-ED [74], RankED [60], and EdgeSAM [76].
@@
Table 4 illustrates that our LUS-Net achieves strong performance across different image modalities (RGB, HHA, and RGB-HHA) and evaluation protocols (S-Eval and C-Eval). For RGB input under S-Eval, although our method achieves slightly lower scores (ODS=0.757) compared to recent methods like RankED (ODS=0.780) and EdgeSAM (ODS=0.783), it maintains competitive performance while using significantly fewer parameters (70.41M vs. 114.4M for RankED). More importantly, our method shows remarkable robustness in the challenging C-Eval protocol. For RGB input, LUS-Net achieves the ODS score of 0.540, significantly outperforming all recent SOTA methods including EDTER (0.416), RankED (0.477), and GCB-Net (0.456). This substantial performance gap (29.8% higher than EDTER, 13.2% higher than RankED, and 18.4% higher than GCB-Net) demonstrates our method’s enhanced capability in generating crisp edge maps. Similar robustness advantages can also be observed in HHA (0.435) and RGB-HHA (0.513) modalities, where our method consistently maintains superior performance across different image types. Specifically, LUS-Net achieves ODS=0.717 in HHA and ODS=0.768 in RGB-HHA under S-Eval protocol, while maintaining strong robustness in C-Eval with ODS=0.435 in HHA and ODS=0.513 in RGB-HHA modalities. In terms of computational efficiency, LUS-Net achieves 34.26 FPS on an A40 GPU, which is significantly faster than transformer-based methods including EDTER (2.1 FPS) and Boosting-ED (2.1 FPS). The Precision-Recall curves in Fig. 9 show that our LUS-Net has a top performance. These results demonstrate that our method achieves an excellent balance between accuracy, robustness, and efficiency.
@@
BIPED: Finally, we further evaluate our method on the BIPED dataset. Since the resolution of each image is relatively large ($1280\times 720$), we crop each image-label pair randomly into $320\times 320$ for training. We adopt recent SOTA methods for comparison which consists of SED [80], HED [24], RCF [25], CED-ADM [81], BDCN [29], PiDiNet [33], CATS [54], FCL-Net [55], DexiNed [45], EDTER [57], and Li et al. [82]. The quantitative comparison results are listed in Table 5 and these results are similar to the performance shown on the BSDS500 and NYUD-V2. Specifically, our LUS-Net achieves ODS=0.902, OIS=0.908, and AP=0.912, surpassing most existing methods. When employing multi-scale testing, the performance further improves to ODS=0.903, OIS=0.909, and AP=0.930. Notably, LUS-Net is the first method to exceed 0.9, demonstrating its powerful performance. Furthermore, compared to other SOTA methods including FCL-Net, EDTER, and DexiNed, our approach achieves more efficient architectural design while maintaining superior performance. These results further validate the effectiveness and generalization ability of our method across different datasets and scenarios.

### v5

@@ 1 Introduction
Edge detection is a fundamental problem in computer vision and image processing, playing a crucial role in a wide range of applications, from autonomous driving to medical imaging. It serves as the cornerstone for many high-level computer vision tasks by identifying the boundaries and structures within images, such as salient object detection [1, 2], object recognition [3, 4], and image segmentation [5, 6]. With the rapid development of deep learning techniques, numerous researchers aim to employ Deep Convolutional Neural Networks (DCNNs) for edge detection. Recent studies have proposed several remarkable DCNN-based methods, such as HED [7], RCF [8], and DexiNed [9]. These DCNN-based approaches have achieved state-of-the-art performance on a number of benchmarks including BSDS500 [10] and NYUD-V2 [11], which demonstrate the strong power of DCNN.
@@
In this work, we attempt to tackle this central issue by addressing the two subsidiary problems. Our goal is to make the network generate one-pixel width edge lines without using the post-processing operation. On the one hand, most DCNN-based methods [7, 8, 12, 13] leverage the powerful feature extraction capabilities of deep learning for edge detection. However, these methods neglect the utilization of prior knowledge regarding image edges, which are typically determined by abrupt changes in pixels. Image second-order derivative captures such changes and it precisely locates the positions of edge pixels by the zero-crossing points. Therefore, leveraging this characteristic can help DCNNs to locate the true edge pixels more accurately. In addition, the mere introduction of second-order derivative information can result in noisy edge maps, as illustrated in Fig. 1 (c). In response to this circumstance, we combine the second-order derivative information with the multi-scale contextual information, developing a Second-order Derivative-based Multi-scale Contextual Enhancement Module (SDMCM). The SDMCM is based on the traditional Laplacian operator [14] which can provide image second-order derivative cues, and dilated convolution with a large receptive field to capture the long-range multi-scale contextual information which can filter out the noise. The proposed SDMCM empowers the network to attain higher precision in true edge pixel localization, consequently generating crisp and clean edge maps.
@@ 2 Related work
Traditional methods: Early edge detection methods typically calculate image derivatives to produce edge maps. The Sobel operator [22] was another first-order edge detector that used two $3\times 3$ kernels to calculate image gradients in horizontal and vertical directions. The Canny detector [23] was a robust algorithm that detected edges by reducing noise, calculating gradients, and applying non-maximum suppression and hysteresis thresholding. The Laplacian detector [14] located edge pixels by computing the second-order derivative of the image intensity, highlighting regions of rapid intensity change. Over time, some researchers improved edge detection by integrating texture, gradient, and other low-level features. Methods like gPb [10] and SE [24] used a classifier to generate object-level boundaries by utilizing these features. Despite their enhanced performance over derivative-based methods, they still relied on human-designed features and lacked semantic information, limiting further improvement.
@@
Deep learning-based methods: Deep learning-based techniques significantly advanced the area of edge detection and played a crucial role in most state-of-the-art (SOTA) edge detection methods. SOTA edge detectors in recent years have mainly adopted convolutional neural networks. These methods achieved significant performance with higher F-scores and some of them even surpassed humans on benchmarks such as BSDS500 dataset. Xie et al. [7] presented the first end-to-end edge detection architecture named Holistically-nested Edge Detection (HED), which was built on a fully convolutional VGG-16 network. They generated edge maps by fusing five side-output features with different scales and constructed a weighted cross-entropy loss function to address the problem of imbalanced pixel distribution. Based on HED, Liu et al. [8] further utilized multi-scale features by aggregating all features from different convolutional layers in each stage of VGG-16. This improved the network’s ability to capture contextual information, making it the first to outperform humans on BSDS500. He et al. [12] proposed an innovative bi-directional cascade network architecture, which leveraged information flow from shallow-to-deep and deep-to-shallow to capture multi-scale features within images comprehensively, significantly enhancing the performance. Soria et al. [9] proposed DexiNed, a network inspired by HED and Xception, which could produce detailed edge maps that were visually appealing without any pre-training or fine-tuning process. Su et al. [15] provided a lightweight yet effective solution for edge detection by integrating traditional edge detection operators into vanilla convolution in modern DCNN.
@@
Regarding precise edge detection, various methods have contributed significantly to the field. Wang et al. [25] focused on guiding the network to predict sharp edge maps by employing sub-pixel convolution to upsample features. Deng et al. [26] explained the reason for edge thickness and proposed a new loss function based on the Dice coefficient which enables the network to generate crisp boundaries without requiring post-processing. Cao et al. [13] presented an innovative network by stacking refinement modules and implementing an adaptive weighting strategy for the optimal combination between different loss functions, resulting in a competitive performance. Huan et al. [27] presented a context-aware tracing strategy using a novel tracing loss and a context-aware fusion block for crisp edge detection. Xuan et al. [28] proposed a novel network architecture designed to enhance the accuracy of edge detection through fine-scale corrective learning mechanisms. Liu et al. [29] proposed GCB-Net, a U-shape encoder-decoder architecture with feature enhancement and fusion modules, which could generate crisp and clean edge maps through multi-scale features and a novel mixed loss function based on Tversky index. Pu et al. [30] EDTER, the first Transformer-based edge detector with a two-stage architecture. The model could extract well-defined boundaries of objects and meaningful edges by utilizing contextual information from the entire image, as well as detail local cues. Zhou et al. [31] designed an edge granularity network to estimate edge granularity from annotations and integrated it into both spatial and frequency domains of feature maps, enabling the generation of diverse edge predictions ranging from coarse object contours to fine textures. Cetinkaya et al. [32] constructed a ranking-based approach to address both class imbalance and label uncertainty in edge detection through ranking positive pixels over negatives and promoting high confidence edge pixels.
@@
Despite their remarkable performance, existing methods still fall short in adequately addressing the challenge of edge thickness. Our approach is motivated by these pioneering works [26, 16, 29] to address the problem of edge thickness. Through the combination of multi-scale contextual and second-order derivative information, we enable the network to produce crisp and clean edge maps. Additionally, our method successfully addresses the issue of imbalanced pixel distribution by constructing a hybrid focal loss function, resulting in excellent performance.
@@ 4.1 Implementation details
Our network is built using the Pytorch deep learning framework. During the training phase, the hyperparameters are as follows: the mini-batch size is 8, the initial learning rate is $1\times 10^{-4}$, the learning rate decay is 0.1, the weight decay is $5\times 10^{-4}$ and the number of training epochs is 40. We decay the learning rate every 5 epochs and adopt the Adam method for optimization. All the experiments are performed using a single Tesla A40 GPU.
@@ 4.3 Ablation experiments on BSDS500 dataset
Loss function: Firstly, we perform ablation experiments on various loss functions to demonstrate the effectiveness of our HFL. To identify the optimal configuration of HFL, we conduct a series of comparison experiments with different parameters, and the comparison results are shown in rows 1-6 of Table 1. It can be observed that, as $\lambda$ decreases from 1 to 0.001, there is a notable increase in C-Eval metrics, while S-Eval metrics maintain relatively stable performance. Such observation can be explained by the complementary roles of the two loss components: focal Tversky loss serves to regulate the crispness of edge delineation, whereas focal loss ensures the accuracy of pixel-level classification in the predicted edge maps. Therefore, the optimal value of $\lambda$ is 0.001. Additionally, when fixed $\lambda=0.001$, the best results of both S-Eval and C-Eval are obtained at $\beta=0.7$, outperforming other values (0.6 and 0.8). In rows 7-10, it can be observed that the best performance is obtained at $\gamma=0.75$. Furthermore, we compare our HFL loss with other loss functions, including Weighted Cross-Entropy [7] (WCE), Hybrid Quality Focal Loss [35] (HQFL), and Hybrid Varifocal Loss [36] (HVFL). The comparison results shown in rows 11-14 demonstrate that the HFL maintains competitive performance in S-Eval (ODS=0.805, OIS=0.827, AP=0.805). More importantly, HFL demonstrates superior performance across all metrics in C-Eval (ODS=0.698, OIS=0.705, AP=0.754), notably outperforming WCE, HQFL, and HVFL. These results strongly indicate that our proposed HFL effectively enhances the model’s capability to produce crisp and clean edge maps while maintaining robust performance in pixel-level classification.
@@
Compression ratio: We also explore the effect of different channel compression ratios $r$ in SDMCM on the model performance and the experiment results are shown in rows 15-17 of Table 1 and Fig. 7. As the ratio increases, the number of model parameters also increases, resulting in a higher computational cost. However, the performance does not improve with an increasing compression ratio. The best performance is obtained at $r=\frac{1}{4}$. The reason for this phenomenon is that a higher compression ratio causes overfitting, resulting in a lower performance both in S-Eval and C-Eval. Therefore, we set the compression ratio to $\frac{1}{4}$, which achieves a balance between performance and parameters.
@@
Image derivative information: In addition, we investigate the effectiveness of different derivative operators by replacing the Laplacian operator in SDMCM with first-order derivative operators (Sobel and Scharr). As shown in rows 18-20 of Table 1, both alternative operators lead to consistent performance degradation across all metrics, substantiating our theoretical analysis that second-order derivative information provides more discriminative features for edge detection compared to first-order derivatives.
@@ 4.4 Comparison with SOTA methods
BSDS500: Firstly, we compare our method with top-performing algorithms on BSDS500. We select recent state-of-the-art edge detectors which can be divided into two categories: the first category is methods without deep learning, which includes Canny [23], gPb-UCM [10], and SE [24]; the second category is approaches using deep learning technique, which including HED [7], RCF [8], BDCN [12], CED [25], LPCB [26], DRC [13], PiDiNet [15], CATS [27], FCL-Net [28], EDTER [30], Boosting-ED [37], DexiNed [9], BetaNet [38], GCB-Net [29], EdgeSAM [39], MuGE [31], and RankED [32]. We additionally reference the studies of prior researchers who utilize extra training data sourced from the PASCAL VOC Context dataset [40] and adopt multi-scale testing to further improve the performance. Table 3 presents all metrics including S-Eval, C-Eval, Params, and FPS. Fig. 8 shows some qualitative comparison results, and Fig. 9 draws the Precision-Recall curves in S-Eval.
@@
NYUD-V2: Secondly, we select the NYUD-V2 dataset to conduct another set of comparison experiments. We adopt some methods as before, which consists of algorithms without using deep learning such as gPb-UCM [10], gPb+NG [41], SE [24] and SE+NG+ [42], and recent top edge detectors based on deep learning such as HED [7], RCF [8], BDCN [12], CATS [27], DRC [13], PiDiNet [15], DexiNed [9], LPCB [26], BetaNet [38], GCB-Net [29], EDTER [30], Boosting-ED [37], RankED [32], and EdgeSAM [39].
@@
Table 4 illustrates that our LUS-Net achieves strong performance across different image modalities (RGB, HHA, and RGB-HHA) and evaluation protocols (S-Eval and C-Eval). For RGB input under S-Eval, although our method achieves slightly lower scores (ODS=0.757) compared to recent methods like RankED (ODS=0.780) and EdgeSAM (ODS=0.783), it maintains competitive performance while using significantly fewer parameters (70.41M vs. 114.4M for RankED). More importantly, our method shows remarkable robustness in the challenging C-Eval protocol. For RGB input, LUS-Net achieves the ODS score of 0.540, significantly outperforming all recent SOTA methods including EDTER (0.416), RankED (0.477), and GCB-Net (0.456). This substantial performance gap (29.8% higher than EDTER, 13.2% higher than RankED, and 18.4% higher than GCB-Net) demonstrates our method’s enhanced capability in generating crisp edge maps. Similar robustness advantages can also be observed in HHA (0.435) and RGB-HHA (0.513) modalities, where our method consistently maintains superior performance across different image types. Specifically, LUS-Net achieves ODS=0.717 in HHA and ODS=0.768 in RGB-HHA under the S-Eval protocol, while maintaining strong robustness in C-Eval with ODS=0.435 in HHA and ODS=0.513 in RGB-HHA modalities. In terms of computational efficiency, LUS-Net achieves 34.26 FPS on an A40 GPU, which is significantly faster than transformer-based methods including EDTER (2.1 FPS) and Boosting-ED (2.1 FPS). The Precision-Recall curves in Fig. 9 show that our LUS-Net has top performance. These results demonstrate that our method achieves an excellent balance between accuracy, robustness, and efficiency.
BIPED: Finally, we further evaluate our method on the BIPED dataset. Since the resolution of each image is relatively large ($1280\times 720$), we crop each image-label pair randomly into $320\times 320$ for training. We adopt recent SOTA methods for comparison which consist of HED [7], RCF [8], BDCN [12], PiDiNet [15], CATS [27], FCL-Net [28], DexiNed [9], and EDTER [30].
As shown in Fig. 11, our LUS-Net demonstrates superior performance compared to existing SOTA approaches on the BIPED dataset. While HED and RCF demonstrate reasonable edge detection capabilities, they suffer from notable limitations including fragmented edges, inconsistent thickness, and discontinuity issues, particularly in complex architectural structures. PiDiNet occasionally misses fine structural details while DexiNed sometimes introduces spurious edges that deviate from the groundtruth. In contrast, our method demonstrates substantial advantages across multiple evaluation criteria, exhibiting enhanced edge continuity that maintains structural integrity while preserving fine-grained details often lost in competing approaches. Our method achieves superior noise suppression, producing cleaner edge maps with reduced artifacts, and demonstrates significantly improved edge localization accuracy with crisper and more precise boundary delineation that closely matches the groundtruth annotations.
Figure 11: Some examples from SOTA methods on BIPED.
@@
The quantitative comparison results are listed in Table 5 and these results are similar to the performance shown on the BSDS500 and NYUD-V2. Specifically, our LUS-Net achieves ODS=0.902, OIS=0.908, and AP=0.912, surpassing most existing methods. When employing multi-scale testing, the performance further improves to ODS=0.903, OIS=0.909, and AP=0.930. Notably, LUS-Net is the first method to exceed 0.9, demonstrating its powerful performance. Furthermore, compared to other SOTA methods including FCL-Net, EDTER, and DexiNed, our approach achieves more efficient architectural design while maintaining superior performance. These results further validate the effectiveness and generalization ability of our method across different datasets and scenarios.